The area of a rectangle is 48 square centimeters. If the length of the rectangle is 8 centimeters longer than the width, how wide is the rectangle?
Answer by RAY100(1637)

Area = Length * Width
.
but l= w+8,,,,,subst
.
A= (w+8)(w) = w^2 +8w
.
but A=48,,,so 48 = w^2 +8w,,,,,or
.
w^2 +8w -48 =0,,,,factor
.
(w+12)(w-4) = 0
.
w= 4,,,,-12(not practical)
.
and,, L= 8+w = 8+4 =12
.
.
check ,,,A=12 *4 = 48,,,,ok
.